OpenStudy (gawain):

Find the derivative using logarithmic differentiation or implicit differentiation : U=√(t+1)/∛(t-1)

OpenStudy (zzr0ck3r):

where do you need help?

OpenStudy (mww):

Take logs of both sides Recall \[\log(x^n)=n logx\] Then also use the result \[\ln(y) = f(x) \rightarrow \frac{ d }{ dx } \ln(y) = \frac{ d }{ dy }\ln(y)\frac{ dy }{ dx } = \frac{ d }{ dx }f(x) \] \[\frac{ d }{ dy }\ln(y) \frac{ dy }{ dx } = \frac{ 1 }{ y } \frac{ dy }{ dx } = f'(x) \rightarrow \frac{ dy }{ dx } = yf'(x)\]